OpenStudy (anonymous):

Mrs. Jones had 13 pairs of old shoes. She went to the store and bought 5 new pairs of shoes. Then she threw away 6 pairs of her old shoes. How many pairs of shoes does she have now?

OpenStudy (anonymous):

13+5=18-6=?

OpenStudy (studygurl14):

She had 13 bought 5 threw away 6 13 + 5 - 6 = ?

OpenStudy (countrygirl1431):

12 i think

OpenStudy (anonymous):

no 16

OpenStudy (anonymous):

@StudyGurl14 @countrygirl1431 can you help me?

OpenStudy (anonymous):

nooo 16 is not it

OpenStudy (anonymous):

get off my page

OpenStudy (anonymous):

12.

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

ok

OpenStudy (studygurl14):

12 is correct.

OpenStudy (anonymous):

^ You're welcome :P

OpenStudy (countrygirl1431):

:p :p :p